# Ctrl-Adapter: An Efficient and Versatile Framework for Adapting Diverse Controls to Any Diffusion Model

### v1

@@ 3.3 Multi-Condition Generation via Ctrl-Adapter Composition
For more effective spatial control beyond a single condition, we can easily combine the control features of multiple ControlNets via our Ctrl-Adapter. As shown in Fig. 4, we first average the ControlNet output features via learnable scalar weights (i.e., $w_{1},w_{2},\dots w_{N}$, $\sum_{i=1}^{N}w_{i}=1$), for each Ctrl-Adapter block, then provide such fused ControlNet features as input to a shared and unified Ctrl-Adapter. The weighted-average approach can be understood as a lightweight mixture-of-experts (MoE) [58]. For training the weights, we first randomly select an integer $k$ from the set $\{1,2,...,N_{\text{max}}\}$, where $N_{\text{max}}$ is a hyper-parameter that balances training speed and composition (default as 4 in our experiments). Then we randomly choose $k$ ControlNets and average the outputs of the selected ControlNet with the learned weights.
@@ 4 Experimental Setup
4.1 ControlNets and Target Diffusion Models
ControlNets.
We use ControlNets trained with SDv1.5.222https://huggingface.co/lllyasviel/ControlNet SDv1.5 has the most number of publicly released ControlNets and has a much smaller training cost compared to recent image/video diffusion models. Note that unlike X-Adapter [48], Ctrl-Adapter does not need to load the source diffusion model (SDv1.5) during training or inference (see (a) and (d) in Fig. 17 for model architecture comparison).
Target diffusion models (where ControlNets are to be adapted).
For video generation models, we experiment with one text-to-video generation model – Hotshot-XL [41], and two image-to-video generation models – I2VGen-XL [75] and Stable Video Diffusion (SVD) [3]. For image generation model, we experiment with the base model in SDXL [44]. For all models, we use their default settings during training and inference (e.g., number of output frames, resolution, number of denoising steps, classifier-free guidance scale, etc.).
4.2 Training Datasets for Ctrl-Adapter
Video datasets.
For training Ctrl-Adapter for video diffusion models, we use 0.2M videos randomly sampled from the Panda-70M training set [7]. Following recent works [3, 9], we filter out videos of static scenes by removing videos whose average optical flow [11, 4] magnitude is below a certain threshold. Concretely, we use the Gunnar Farneback’s algorithm333 https://docs.opencv.org/4.x/d4/dee/tutorial_optical_flow.html [12] at 2FPS, calculate the averaged the optical flow for each video and re-scale it between 0 and 1, and filter out videos whose average optical flow error is below a threshold of 0.25. This process gives us a total of 0.2M remaining videos.
Image datasets.
4.3 Input Conditions
4.4 Evaluation datasets
Video datasets.
Image datasets.
4.5 Evaluation Metrics
Visual quality.
Spatial control.
Table 2: Evaluation of video generation with single control condition (depth map/canny edge) on DAVIS 2017 dataset. : skipping latents $\bm{z}$ to ControlNet input to handle different noise scale (see Sec. 3.2 for details).
@@ 4.5 Evaluation Metrics
SDv1.5 ControlNet + Ctrl-Adapter (Ours) 4.05 3.68 4.16 3.96
@@ 5 Results and Analysis
We demonstrate the usefulness of Ctrl-Adapter with comprehensive experiments and analysis. We first evaluate Ctrl-Adapter for controlled video (Sec. 5.1) and image generation (Sec. 5.2). Then we evaluate the combination of multiple Ctrl-Adapter for video generation with multiple conditions (Sec. 5.3). We also experiment video generation with sparse frame conditions (Sec. 5.4) and zero-shot generalization on unseen conditioned modalities (Sec. 5.5). Moreover, we analyze the training efficiency of Ctrl-Adapter (Sec. 5.6). Lastly, we showcase a video editing pipeline by controlling both image and video diffusion models with Ctrl-Adapter (Sec. 5.7).
5.1 Video Generation with Single Control Condition
Figure 5: Video generation with different video control methods and Ctrl-Adapter using depth map (left) and canny edge (right) conditions. See full videos on our project page.
Figure 6: Videos generated from different video control methods and Ctrl-Adapter on DAVIS 2017, using depth map (left) and canny edge (right) conditions. See full videos on our project page.
We experiment with Ctrl-Adapter on a text-to-video (T2V) generation backbone, Hotshot-XL [41], and two image-to-video (I2V) generation backbones, I2VGen-XL [75] and Stable Video Diffusion (SVD) [3]. We compare our method with previous controllable video generation models, including Text2Video-Zero [28], Control-A-Video [8], ControlVideo [76], and VideoComposer [66]. As the spatial layers of Hotshot-XL are initialized with SDXL and remain frozen, the SDXL ControlNets are directly compatible with Hotshot-XL, so we include Hotshot-XL + SDXL ControlNet as a baseline. We also experiment with a temporal ControlNet [53] trained with SVD. While previous works (ControlVideo, Control-A-Video, Text2Video-Zero) generate 8-frame videos, Ctrl-Adapter can flexibly adapt ControlNets to different video backbones that output equal length or even longer videos: Hotshot-XL (8 frames), SVD (14 frames) and I2VGen-XL (16 frames). Following Hu et al[27], we evaluate all methods with FID and optical flow error.
SDv1.5 ControlNet + Ctrl-Adapter outperforms other video control methods.
Table 2 shows that in both depth map and canny edge input conditions, Ctrl-Adapters on I2VGen-XL and SVD outperforms all previous strong video control methods in visual quality (FID) and spatial control (optical flow error) metrics. Note that it takes less than 10 GPU hours for Ctrl-Adapter to outperform the baselines (see Sec. 5.6). In Fig. 5, we compare Ctrl-Adapter and other recent video control methods. When applied to three different backbones, Ctrl-Adapter can handle challenging big/rapid movement controls. In Fig. 6, we also show videos generated on DAVIS 2017 dataset (described in Sec. 4.4).
Table 3: Evaluation of image generation with single control condition (depth map/canny edge) on COCO val2017 split.
Method Depth Map Canny Edge
FID ($\downarrow$) MSE ($\downarrow$) SSIM ($\uparrow$) FID ($\downarrow$) SSIM ($\uparrow$)
@@ 5.1 Video Generation with Single Control Condition
SDv1.5 ControlNet [74] 21.25 87.57 - 18.90 0.4828
T2I-Adapter [40] 21.35 89.82 - 18.98 0.4422
GLIGEN [33] 21.46 88.22 - 24.74 0.4226
Uni-ControlNet [78] 21.20 91.05 - 17.79 0.4911
@@
SDXL ControlNet [62] 17.91 86.95 0.8363 17.21 0.4458
SDv1.5 ControlNet + X-Adapter [48] 20.71 90.08 0.7885 19.71 0.3002
SDv1.5 ControlNet + Ctrl-Adapter (Ours) 18.68 87.54 0.8515 21.77 0.5101
Figure 7: Image generation results on COCO val2017 split using depth map and canny edge conditions. Ctrl-Adapter achieves better spatial control compared to X-Adapter (e.g., the arm of the metal statue for the second depth map input, the location of the person for the last depth map input, the tilt of the kitten’s head for the first canny edge input, and the bike’s orientation for the last canny edge input) and matches the performance of SDXL ControlNet.
5.2 Image Generation with Single Control Condition
@@
We compare Ctrl-Adapter with recent image control methods that use SDv1.4, SDv1.5 and SDXL as backbones. For methods using SDv1.4/v1.5 as a backbone, we compare Ctrl-Adapter to pretrained SDv1.5 ControlNet [74], GLIGEN [33], T2I-Adapter [40], and Uni-ControlNet [78]. For methods using SDXL as a backbone, we compare Ctrl-Adapter to pretrained SDXL ControlNet [62] and X-Adapter [48]. Following Zhao et al[78], we report the FID, MSE, and SSIM scores.
SDv1.5 ControlNet + Ctrl-Adapter outperforms other baselines and matches SDXL ControlNet.
@@
As shown in Table 3, we find that Ctrl-Adapter outperforms all baselines with SDv1.4/v1.5 in most metrics (FID/MSE with depth map inputs and SSIM with canny edge inputs). When compared to the baselines with SDXL backbones, Ctrl-Adapter outperforms X-Adapter in most metrics, and matches (in FID/MSE with depth map inputs) or outperforms SDXL ControlNet (in SSIM with depth map and canny edge inputs). Note that SDXL ControlNet was trained for much longer than Ctrl-Adapter (700 vs. 22 A100 GPU hours) and it takes less than 10 GPU hours for Ctrl-Adapter to outperform the SDXL ControlNet in SSIM in depth map (see Sec. 5.6). In Fig. 7, we show the image control methods with SDXL backbones. Ctrl-Adapter achieves better spatial control compared to X-Adapter (e.g., the arm of the metal statue for the second depth map input, the location of the person for the last depth map input, the bear’s orientation for the third canny edge input, and the bike’s orientation for the last canny edge input) and matches the performance of SDXL ControlNet.
@@
Table 4: Comparison of single and multi-condition video generation with Ctrl-Adapter. We adapt SDv1.5 ControlNets to I2VGen-XL and use DAVIS 2017 dataset for evaluation. The control sources are abbreviated as D (depth map), C (canny edge), N (surface normal), S (softedge), Seg (semantic segmentation map), L (line art), and P (human pose).
Conditions FID ($\downarrow$) Optical Flow Error ($\downarrow$)
Single condition
D 7.43 3.20
C 6.42 3.37
Multiple conditions with equal weights
D + C 8.50 2.84
D + P 11.32 3.48
D + C + N + S 8.75 2.40
D + C + N + S + Seg + L + P 9.48 2.93
Multiple conditions with linear weights (default)
D + C 9.14 2.89
D + P 10.98 3.32
D + C + N + S 8.39 2.23
D + C + N + S + Seg + L + P 8.18 2.48
Figure 8: Video generation from single and multiple conditions with Ctrl-Adapter on I2VGen-XL. (a) single condition: depth map; (b) 2 conditions: depth map + human pose; (c) 4 conditions: depth map + human pose + canny edge + semantic segmentation. While all videos correctly capture high-level dynamics, adding more conditions can help fix several minor artifacts (e.g., in (a) – building is blurred; in (b) – purse color changes).
Figure 9: Additional video generation from 4 conditions: depth map + human pose + canny edge + semantic segmentation. See full videos on our project page.
@@ 5.3 Video Generation with Multiple Control Conditions
As described in Sec. 3.3, users can easily achieve multi-source control by simply combining the control features of multiple ControlNets via our Ctrl-Adapter. In Table 4, we compare the multi-source and single-source controlled video generation results with I2VGen-XL backbone in terms of visual quality (FID) and spatial control (optical flow error).
More conditions improve spatial control in video generation.
Table 4 shows a clear trend that control with multiple conditions almost always yields better spatial control (lower optical flow error) and slightly lower visual quality (higher FID) than control with a single condition. See also Sec. D.2 for additional study on visual quality-spatial control tradeoff. Among different conditions, we find the combination of depth map and canny edges is more helpful than depth map and human pose, perhaps because not all videos include humans. Fig. 8 shows example videos generated with single and multiple conditions. While all videos correctly capture the high-level dynamics of ‘a woman wearing purple strolling during sunset’, the videos generated with more conditions show more robustness in several minor artifacts. For example, when only depth map is given (Fig. 8 a), the building behind the person is blurred. When depth map and human pose are given (Fig. 8 b), the color of the purse changes from white to purple. When four conditions (depth map, canny edge, human pose, and semantic segmentation) are given, such artifacts are removed (Fig. 8 c). Fig. 9 shows additional generation examples with the four conditions (a woman with a child / a dancing man).
Linear weight is better than equal weights.
When comparing equal weights and linear weights for averaging ControlNet features, Table 4 shows that in both FID and optical flow error, the linear weight outperforms the equal-weight approach as the number of conditions increases. This performance gain arises from the flexibility that allows various ControlNet experts to contribute differently in each middle and output block. We visualize the learned weights for each control condition across all mid/output blocks in Fig. 25. In Table 6, we show additional experiments with learning a small MLP-based router module to assign weights to fuse ControlNet outputs, where the simple linear weights perform similarly.
Figure 10: Video generation from sparse frame conditions with Ctrl-Adapter on I2VGen-XL backbone. I2VGen-XL generates 16-frame videos, where we provide sparse control conditions (user scribble or depth map) for 1st, 6th, 11th, and 16th frames. See full videos on our project page.
@@ 5.4 Video Generation with Sparse Frames as Control Condition
We experiment Ctrl-Adapter with providing sparse frame conditions (i.e., conditions are given for only a subset of frames), using the I2VGen-XL backbone. During each training step, we first randomly select an integer $k\in\{1,...,N\}$, where $N$ is equal to the total number of output frames of the video generation backbone (e.g., $N=16$ for I2VGen-XL). Next, we randomly select $k$ keyframes from $N$ total frames. We then extract these keyframes’ depth maps and user scribbles as control conditions. As explained in Sec. 3.2, we do not give the latents $\bm{z}$ and only give the $k$ frames to ControlNet.
Ctrl-Adapter can effectively handle sparse video control conditions.
In Fig. 10, we can see that I2VGen-XL with our Ctrl-Adapter can correctly generate videos that follow the scribble conditions for the given 4 sparse keyframes (out of a total of 16) and make reasonable interpolations on the frames without conditions. In Sec. 6.2, we also show that skipping the latent from ControlNet inputs is important in improving the sparse control capability.
@@ 5.5 Zero-Shot Generalization on Unseen Conditions
Since the majority of ControlNet parameters are the same as their backbone generation models, ControlNet can be interpreted as an image feature extractor that maps different types of images to the unified representation of backbone generation models. This begs an interesting question: “Does Ctrl-Adapter learn general feature mapping from one (smaller) backbone to another (larger) backbone model?” To answer this question, we experiment with directly using Ctrl-Adapter to adapt ControlNets that Ctrl-Adapter have not seen during training. In Fig. 11, we show the Ctrl-Adapter trained with a depth map SDv1.5 ControlNet can successfully adapt different SDv1.5 ControlNets that encode surface normal map, line art, and softedge conditions in a zero-shot manner.

### v2

@@ 3.3 Multi-Condition Generation via Ctrl-Adapter Composition
Multi-ControlNet [86] is proposed for spatial control beyond a single condition. However, this method naively combines different conditions with equal weights during inference time without training. For more effective control composition, we first experiment with some simple extensions, such as replacing these fixed weights with unconditional global learnable weights via a lightweight MoE [68] router (see variant (a) in Fig. 5 right). Next, we further propose refining the global weights MoE router into a more fine-grained, patch-level MoE router. Variant (b) processes the patch-level features of each ControlNet into a scalar value independently, then uses softmax to assign ControlNet weights for each patch. Variant (c) takes in all $N$ ControlNet features associated with a patch, using an architecture design inspired by Q-Former [37] to output expert weights. Comparisons of different variants are discussed in Sec. 5.3 and Sec. E.4.
@@ 4 Experimental Setup
ControlNets and Target Diffusion Models. We use ControlNets trained with SD v1.5. For target diffusion models, we experiment with two I2V models – I2VGen-XL [89] and Stable Video Diffusion (SVD) [3], two T2V models – Latte [46] and Hotshot-XL [49], and two T2I models – SDXL [52] and PixArt-$\alpha$ [8]. Note that Latte and PixArt-$\alpha$ are generation models based on DiT instead of U-Net.
Training and Evaluation Datasets. We use 200K videos sampled from Panda-70M training set [9] and 300K images from the LAION POP [67] dataset for video and image Ctrl-Adapters training respectively. During training, we extract various control conditions (e.g., depth map) on-the-fly to simplify the data-preparation process. Following previous works [31, 90], we evaluate video Ctrl-Adapters on DAVIS 2017 [53], and image Ctrl-Adapters on COCO val2017 split [42]. Detailed training and inference setups for the experiments are provided in Appendix C and Appendix D.
Evaluation Metrics. We perform evaluation on two folds: visual quality and spatial control. Following previous works [54, 31], we use Frechet Inception Distance (FID) [26] to measure the visual quality of generated images/videos. For video datasets, following VideoControlNet [31], we report the L2 distance between the optical flow error [58] between the input and generated videos. For image datasets, following Uni-ControlNet [92], we report the Structural Similarity (SSIM) [78] and mean squared error (MSE) between generated images and ground truth images.
Table 2: Evaluation of video generation with single control condition on DAVIS 2017 dataset. The best number in each column is bolded, and the second best is underscored.
@@
Latte backbone (DiT-Based)
SDv1.5 ControlNet + Ctrl-Adapter (Ours) 16.92 3.98 17.87 2.73
@@
SDv1.5 ControlNet + Ctrl-Adapter (Ours) 3.82 2.96 3.96 2.39
@@ 5 Results and Analysis
5.1 Video Generation with Single Condition
Figure 6: Single-cond. video generation.
We compare SDv1.5 ControlNet + Ctrl-Adapter built on Hotshot-XL, I2VGen-XL, SVD, and Latte with video control methods including Text2Video-Zero [32], Control-A-Video [10], ControlVideo [90], and VideoComposer [77]. As the spatial layers of Hotshot-XL are initialized with SDXL and remain frozen, the SDXL ControlNets are directly compatible with Hotshot-XL, so we include Hotshot-XL + SDXL ControlNet as a baseline. We also experiment with a temporal ControlNet [62] trained with SVD.
Table 2 shows that in both depth map and canny edge input conditions, Ctrl-Adapters on I2VGen-XL and SVD outperforms all previous strong video control methods in visual quality (FID) and spatial control (optical flow error) metrics. Note that it takes $<$ 10 GPU hours for Ctrl-Adapter to outperform the baselines (see Fig. 2). In Fig. 6 and Sec. H.1, we visualize the comparison between Ctrl-Adapter and other video control baselines. We study visual quality-spatial control trade-off in Sec. F.2.
Table 3: Evaluation of image generation with single control condition on COCO val2017 split. The best number in each column is bolded, and the second best is underscored.
Method Depth Map Canny Edge Soft Edge / HED
FID ($\downarrow$) MSE ($\downarrow$) SSIM ($\uparrow$) FID ($\downarrow$) SSIM ($\uparrow$) FID ($\downarrow$) SSIM ($\uparrow$)
@@ 5.1 Video Generation with Single Condition
SDv1.5 ControlNet [86] 21.25 87.57 - 18.90 0.4828 26.59 0.4719
T2I-Adapter [48] 21.35 89.82 - 18.98 0.4422 - -
GLIGEN [39] 21.46 88.22 - 24.74 0.4226 28.57 0.4015
Uni-ControlNet [92] 21.20 91.05 - 17.79 0.4911 17.86 0.5197
@@
SDXL ControlNet [73] 17.91 86.95 0.8363 17.21 0.4458 - -
SDv1.5 ControlNet + X-Adapter [56] 20.71 90.08 0.7885 19.71 0.3002 - -
SDv1.5 ControlNet + Ctrl-Adapter (Ours) 19.26 87.54 0.8534 21.04 0.5806 18.08 0.6454
PixArt-$\alpha$ backbone (DiT-Based)
PixArt-$\delta$ ControlNet [7] - - - - - 20.41 0.6938
SDv1.5 ControlNet + Ctrl-Adapter (Ours) 22.54 84.78 0.8496 18.75 0.6359 17.52 0.6812
Figure 7: Generated images on COCO val2017 split.
5.2 Image Generation with Single Condition
@@
We compare SDv1.5 ControlNet + Ctrl-Adapter with controllable image generation methods that use SDv1.4, SDv1.5, SDXL, and PixArt-$\alpha$ as backbones, including pretrained SDv1.5/SDXL ControlNets [86, 73], T2I-Adapter [48], GLIGEN [39], Uni-ControlNet [92], X-Adapter [56], and PixArt-$\delta$ ControlNet [7].
@@
As shown in Table 3, Ctrl-Adapter outperforms baselines with SDv1.4/v1.5 backbones in almost all metrics. When compared to the baselines with SDXL backbones, Ctrl-Adapter outperforms X-Adapter in most metrics, and matches (in FID/MSE with depth map inputs) or outperforms SDXL ControlNet (in SSIM with depth map and canny edge inputs). Note that SDXL ControlNet was trained for much longer than Ctrl-Adapter (700 vs. 44 A100 GPU hours) and it takes less than 10 GPU hours for Ctrl-Adapter to outperform the SDXL depth ControlNet in SSIM (see Fig. 2). In addition, when applied to DiT-based backbone (i.e., PixArt-$\alpha$), Ctrl-Adapter achieves good improvement in FID (17.52 ours vs. 20.41 PixArt-$\delta$ ControlNet on soft edge) and competitive SSIM score. In Fig. 6 and Fig. 7, we visualize the comparison between Ctrl-Adapter and other image control baselines. See Sec. H.3 for more visualizations.
@@
Table 4: Comparison of different weighting methods (see Fig. 5 right part for details) for multi-condition video generation. The control sources are abbreviated as D (depth map), C (canny edge), N (surface normal), S (softedge), Seg (semantic segmentation map), L (line art), and P (human pose).
D+C D+P D+C+N+S D+C+N+S+Seg+L+P
FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$)
Baseline: Equal Weights 8.50 2.84 11.32 3.48 8.75 2.40 9.48 2.93
(a) Unconditional Global Weights 9.14 2.89 10.98 3.32 8.39 2.36 8.18 2.48
(b) Patch-Level MLP Weights 8.40 2.34 9.37 3.17 7.87 2.11 8.26 2.00
(c) Patch-Level Q-Former Weights 7.54 2.39 9.22 3.22 7.72 2.31 8.00 2.08
@@ 5.3 Video Generation with Multiple Control Conditions
As described in Sec. 3.3, users can achieve multi-source control by simply combining the control features of multiple ControlNets via our Ctrl-Adapter. Table 4 shows the result in two folds: firstly, patch-level MoE routers (i.e., variants b and c in Fig. 5) consistently outperforms the equal weights baseline as well as the unconditional global weights (i.e., variant a in Fig. 5), which proves the effectiveness of patch-level fine-grained control composition. Secondly, as shown in (b) and (c), control with more conditions almost always yields better spatial control and visual quality than control with a single condition. Fig. 28 and Fig. 29 show that multi-condition composition provides more accurate control compared to a single condition. Table 8 extends (a) by conditioning on image/text/timestep embeddings.
Figure 8: Video generation from sparse frame conditions with Ctrl-Adapter on I2VGen-XL (which generates 16 frames in total). We only provide controls for the 1st, 6th, 11th, and 16th frames.
@@ 5.4 Video Generation with Sparse Frames as Control Condition
We experiment Ctrl-Adapter with providing sparse frame conditions using I2VGen-XL as backbone. During each training step, we first randomly select an integer $k\in\{1,...,N\}$, where $N$ is equal to the total number of output frames (e.g., $N=16$ for I2VGen-XL). Next, we randomly select $k$ key frames from $N$ total frames. We then extract these key frames’ depth maps and user scribbles as control conditions. we do not give the latents $\bm{z}$ and only give the $k$ frames to ControlNet. In Fig. 8, we can see that I2VGen-XL with our Ctrl-Adapter can correctly generate videos that follow the control conditions for the given 4 sparse key frames and make reasonable interpolations on the frames without conditions. In Sec. E.3, we show that skipping the latent from ControlNet inputs is important in improving the sparse control capability.
@@ 5.5 Zero-Shot Generalization on Unseen Conditions
ControlNet can be understood as an image feature extractor that maps different types of controls to the unified representation space of backbone generation models. This begs an interesting question: “Does Ctrl-Adapter learn general feature mapping from one (smaller) backbone to another (larger) backbone?” To answer this question, we experiment by directly plugging Ctrl-Adapter to ControlNets that are not seen during training. In Fig. 9, we observe the Ctrl-Adapter trained on depth maps can adapt to normal map and soft edge ControlNets in a zero-shot manner. Quantitative analysis of different training strategies based on such observation is illustrated in Sec. F.1.